Examine the series for convergence for summation of ((-1)^(n-1) sin(nx))/n^3

1
Expert's answer
2022-05-09T12:27:45-0400

"\begin{array}{l}
\text { 1) } \sum_{n=1}^{\infty} \frac{(-1)^{(n-1)} \cdot \sin (n x)}{n^{3}} \quad \text { if } \frac{\sin (n x)}{n^{3}} \rightarrow \text { monotone } \\
\text { for any } x \in N \quad \lim _{n \rightarrow \infty} \frac{\sin (n x)}{n^{3}}=0 \\
\text { So series converges }\\
\end{array}"